OpenStudy (anonymous):

Explain, using wave mechanics, why it is impossible for a particle to have zero kinetic energy when it is confined to a fixed region in space.

OpenStudy (anonymous):

I'm assuming this coming straight from quantum mechanics. If we confine a particle to a well (I shall not derivate QM here), we argue that the potential energy is 0. Therefore, the total energy is all kinetic energy. For a infinite potential well, this energy is given by:\[E_{n}=\frac{n^{2}\pi^{2}\hbar^{2}}{2mL^{2}}\] For the lowest energy possible, n=1:\[E_{1}=\frac{\pi^{2}\hbar^{2}}{2mL^{2}} \neq0\]